### v3

@@ Abstract
Keywords:
Malware AnalysisData PreprocessingLarge Language ModelCybersecurity.
@@ 1.2 Preprocessing and Features Description
Based on our malware analysis experience, on expert knowledge, and inspired by work as EMBER [2], we propose the following preprocessing to convert a binary file into a convenient form to use it with machine learning or deep learning models. We extract substantial information from the file, generally used for static malware detection and analysis, and format this information into a report as a JSON file. As EMBER, we mainly focus on extracting static features contained in binary files, because they are the main characteristics used by malware analysts in this field. However, we also extract information based on rules or signatures to enrich our JSON reports. These rules or signatures give behavioral (MITRE ATT&CK and MBC) or obfuscation (packing) information. In the remainder of this section, we will describe in more details the characteristics extracted from the binary files. For each subsection of the JSON report, we will provide examples based on a variant of the Wannacry ransomware [17], whose sha256 is ed01ebfbc9eb5bbea545af4d01bf5f1071661840480439c6e5babe8e080e41aa.
1.2.1 Global information
The first section of a JSON report provides global information on the analysis file as the file name, signature (sha256, md5) the file type, the target OS, the compilation date, the file size and the file entropy. This information describes the file, and for instance a weird compilation date or a hight entropy can be indicators of a suspicious behavior. However, it cannot be used to decide whether a file is a benign file or a malware, and it needs to be completed with all the information explained below.
@@ 1.2.1 Global information
1.2.2 Section information
The second part of the report gives information related to the sections of the file.
In PE files, the section table describes how the program is organized in memory. Each section contains data or code needed during runtime or linking. A common section includes .text (executable code), .data (initialized data), .rdata (read-only data) and .rsrc (resources like icons or dialog). Each section has a set of characteristics, with a flag describing the section’s attributes, such as whether it is executable, readable, writable or contains code, as described in table 2.
@@ 1.2.2 Section information
It is common to base malware analysis on section analysis [21, 29]. For example, changing section names is a common technique used to evade detection or hinder reverse engineering [1, 22]. In the other hand, some section names are a characteristic of evasion tools such as UPX [^1] https://upx.github.io/ that compress the original code and data into custom sections like .upx1 or .upx2. Additionally, malware may assign unusual characteristics to a section, such as making the .rsrc executable, which may indicate a hidden malicious payload in this section. These irregularities, in naming and characteristics, are red flags and often indicate packing, obfuscation or code injection techniques used to bypass antivirus engines or harden reverse engineering.
In our JSON report, as described in figure 3, we keep, among other things, the name, the size, the signature (sha256) and the characteristics of the sections.
@@
1.2.3 Imported functions
The Import Address Table (IAT) is a critical structure within a PE (Portable Executable) file that lists all the external functions a program relies on, typically from Windows DLLs like kernel32.dll, user32.dll, or advapi32.dll. Each entry in the IAT corresponds to a function that will be resolved and loaded by the Windows loader at runtime, pointing to the actual memory address of the function in the relevant DLL. The IAT contains both the names of the imported functions and the DLLs they come from. During static analysis, the IAT provides valuable insight into a binary’s behavior. For instance, imports by ordinal instead of by name can be suspicious [3], as it is a common tactic used to evade detection or obfuscate a malicious intent. Analysts also look for malicious patterns in imported functions, such as the presence of API calls often used in code injection or process hollowing, for example. A cluster of the imports contained can strongly suggest the binary may be performing one of the actions described in Table 3. Therefore, scrutinizing the IAT is a key step in identifying potential threats during static malware analysis.
@@ 1.2.3 Imported functions
In the next section, we extract information related to the IAT. As described in Figure 4 We first give the imphash, a file signature related to the imported functions, and the number of clear and ordinal imported functions. Then we give the list of all imported functions, sorted by library (kernel32.dll for example). If a function is imported by ordinal, we retrieve, if possible, its name.
Figure 4: IAT information
1.2.4 Packing signatures
The fourth section, illustrated in Figure 5 is related to the obfuscation, in particular if the file is packed. Packing code is a common technique used by threat actor and malware editors to hide malicious payload and disable or prevent static analysis [27]. Even if it is a technique used by software editors for intellectual property reasons, it is also a indicator of suspicious for malicious files analysis.
@@ 1.2.4 Packing signatures
where $l_{pi}\in\{0,1\}$ is the packing label returned by each packing detection tools, with $l_{pi}=1$ if the tools detect a packer signature or indication of packing. Each sub-label $l_{pi}$ is weighted by a weight $w_{i}$ defined by the efficiency of the tools for packing detection.
Finally, based on the signature returned by packing detection tools, we also return the potential packing software used on the file.
@@
1.2.5 Mitre Att&ck and MBC information
Then the last section is related to Mitre Att&ck and MBC identification.
The MITRE ATT&CK [^5] https://attack.mitre.org/ framework is a globally accessible knowledge base of adversary tactics and techniques based on real-world observations. It provides a structured approach to understanding how cyberthreats operate across the attack lifecycle. For malware analysis, especially in categorizing threats like trojans and ransomware, ATT&CK helps analysts to map malicious behaviors to specific techniques such as credential dumping, command and control, or data encryption for impact. By aligning observed malware behaviors with ATT&CK techniques, security professionals can gain deeper insight into the malware’s functionality, objectives, and potential mitigation. This behavioral mapping is critical for threat intelligence, detection engineering, and incident response, allowing defenders to anticipate attacker moves and improve their defensive posture.
@@ 2.2 BERT Results
0 0.96 0.95 0.96 23600
1 0.95 0.97 0.96 13122
2 0.98 0.95 0.97 663
3 0.94 0.99 0.96 5811
4 0.92 0.93 0.93 347
5 0.91 0.93 0.92 845
6 0.91 0.23 0.37 560
7 0.97 0.94 0.96 157
@@
0 0.95 0.94 0.94 5793
1 0.94 0.96 0.95 3392
2 0.94 0.89 0.91 150
3 0.93 0.98 0.95 1471
4 0.88 0.79 0.83 100
5 0.85 0.86 0.86 183
6 0.83 0.29 0.43 153
7 1.00 0.97 0.99 35
@@ 2.3 ModernBERT results
Using the ModernBERT model to have a maximum sequence length from $512$ to $1024$ does not noticeably affect training speed, but it does require more memory. Nevertheless, this trade-off appears justified given the improved malware classification results.
@@
0 0.98 0.96 0.97 23600
1 0.96 0.98 0.97 13122
2 0.99 0.95 0.97 663
3 0.94 0.99 0.97 5811
4 0.95 0.94 0.94 347
5 0.96 0.96 0.96 845
6 0.87 0.82 0.84 560
7 0.99 0.97 0.98 157
@@
0 0.95 0.94 0.94 5793
1 0.94 0.95 0.95 3392
2 0.93 0.89 0.91 150
3 0.92 0.97 0.95 1471
4 0.89 0.78 0.83 100
5 0.86 0.88 0.87 183
6 0.77 0.75 0.75 153
7 0.86 0.89 0.87 35
@@
Overall, these results highlight the potential of leveraging language models associated with very complete reports on malware for automated classification. It is thus important to create such reports, gathering as much information as possible, from static and behavioral analysis.

### v4

@@ 1.2 Preprocessing and Features Description
Based on our malware analysis experience, on expert knowledge, and inspired by work as EMBER [2], we propose the following preprocessing to convert a binary file into a convenient form to use it with machine learning or deep learning models. We extract substantial information from the file, generally used for static malware detection and analysis, and format this information into a report as a JSON file. As EMBER, we mainly focus on extracting static features contained in binary files, because they are the main characteristics used by malware analysts in this field. However, we also extract information based on rules or signatures to enrich our JSON reports. These rules or signatures give behavioral (MITRE ATT&CK and MBC) or obfuscation (packing) information. In the remainder of this section, we will describe in more details the characteristics extracted from the binary files. For each subsection of the JSON report, we will provide examples based on a variant of the Wannacry ransomware [18].
The first section of a JSON report provides global information on the analysis file as the file name, signature (sha256, md5) the file type, the target OS, the compilation date, the file size and the file entropy. This information describes the file, and for instance a weird compilation date or a hight entropy can be indicators of a suspicious behavior. However, it cannot be used to decide whether a file is a benign file or a malware, and it needs to be completed with all the information explained below. Figure 2 provides an example if such a JSON section.
@@
The second part of the report gives information related to the sections of the file. In PE files, the section table describes how the program is organized in memory. Each section contains data or code needed during runtime or linking. A common section includes .text (executable code), .data (initialized data), .rdata (read-only data) and .rsrc (resources like icons or dialog). Each section has a set of characteristics, with a flag describing the section’s attributes, such as whether it is executable, readable, writable or contains code, as described in Table 2.
@@
It is common to base malware analysis on section analysis [22, 31]. For example, changing section names is a common technique used to evade detection or hinder reverse engineering [1, 24]. In the other hand, some section names are a characteristic of evasion tools such as UPX [^1] https://upx.github.io/ that compress the original code and data into custom sections like .upx1 or .upx2. Additionally, malware may assign unusual characteristics to a section, such as making the .rsrc executable, which may indicate a hidden malicious payload in this section. These irregularities, in naming and characteristics, are red flags and often indicate packing, obfuscation or code injection techniques used to bypass antivirus engines or harden reverse engineering. In our JSON report, as described in Figure 3, we keep, among other things, the name, the size, the signature (sha256) and the characteristics of the sections.
@@
The Import Address Table (IAT) is a critical structure within a PE (Portable Executable) file that lists all the external functions a program relies on, typically from Windows DLLs like kernel32.dll, user32.dll, or advapi32.dll. Each entry in the IAT corresponds to a function that will be resolved and loaded by the Windows loader at runtime, pointing to the actual memory address of the function in the relevant DLL. The IAT contains both the names of the imported functions and the DLLs they come from. In our preprocessing, we extract information related to the IAT. As described in Figure 4, we first give the imphash, a file signature related to the imported functions, and the number of clear and ordinal imported functions. Then we give the list of all imported functions, sorted by library (kernel32.dll for example). If a function is imported by ordinal, we retrieve, if possible, its name.
Figure 4: IAT information
During static analysis, the IAT provides valuable insight into a binary’s behavior. For instance, imports by ordinal instead of by name can be suspicious [3], as it is a common tactic used to evade detection or obfuscate a malicious intent. Analysts also look for malicious patterns in imported functions, such as the presence of API calls often used in code injection or process hollowing, for example. A cluster of the imports contained can strongly suggest the binary may be performing one of the actions described in Table 3. Therefore, scrutinizing the IAT is a key step in identifying potential threats during static malware analysis.
@@
The fourth section, illustrated in Figure 5 is related to the obfuscation signatures, in particular if the file is packed. Packing code is a common technique used by threat actor and malware editors to hide malicious payload and disable or prevent static analysis [29]. Even if it is a technique used by software editors for intellectual property reasons, it is also a indicator of suspicious for malicious file analysis.
@@
where $l_{pi}\in\{0,1\}$ is the packing label returned by each packing detection tool, with $l_{pi}=1$ if the tool detects a packer signature or indication of packing. The weights $w_{i}$ are defined by the efficiency of the packing detectors. Based on the final signature , we also return the potential packing software used on the file.
@@
Then the last section is related to Mitre Att&ck and MBC identification. The MITRE ATT&CK [^5] https://attack.mitre.org/ framework is a globally accessible knowledge base of adversary tactics and techniques based on real-world observations. It provides a structured approach to understanding how cyberthreats operate across the attack lifecycle. For malware analysis, especially in categorizing threats like trojans and ransomware, ATT&CK helps analysts to map malicious behaviors to specific techniques such as credential dumping, command and control, or data encryption for impact. By aligning observed malware behaviors with ATT&CK techniques, security professionals can gain deeper insight into the malware’s functionality, objectives, and potential mitigation. This behavioral mapping is critical for threat intelligence, detection engineering, and incident response, allowing defenders to anticipate attacker moves and improve their defensive posture.
@@ 2.2 BERT Results
trojan 0.96 0.95 0.96 23600
worm 0.95 0.97 0.96 13122
ransomware 0.98 0.95 0.97 663
backdoor 0.94 0.99 0.96 5811
infostealer 0.92 0.93 0.93 347
downloader 0.91 0.93 0.92 845
dropper 0.91 0.23 0.37 560
virus 0.97 0.94 0.96 157
@@
trojan 0.95 0.94 0.94 5793
worm 0.94 0.96 0.95 3392
ransomware 0.94 0.89 0.91 150
backdoor 0.93 0.98 0.95 1471
infostealer 0.88 0.79 0.83 100
downloader 0.85 0.86 0.86 183
dropper 0.83 0.29 0.43 153
virus 1.00 0.97 0.99 35
@@ 2.3 ModernBERT results
trojan 0.98 0.96 0.97 23600
worm 0.96 0.98 0.97 13122
ransomware 0.99 0.95 0.97 663
backdoor 0.94 0.99 0.97 5811
infostealer 0.95 0.94 0.94 347
downloader 0.96 0.96 0.96 845
dropper 0.87 0.82 0.84 560
virus 0.99 0.97 0.98 157
@@
trojan 0.95 0.94 0.94 5793
worm 0.94 0.95 0.95 3392
ransomware 0.93 0.89 0.91 150
backdoor 0.92 0.97 0.95 1471
infostealer 0.89 0.78 0.83 100
downloader 0.86 0.88 0.87 183
dropper 0.77 0.75 0.75 153
virus 0.86 0.89 0.87 35
@@
Using the ModernBERT model to have a maximum sequence length from $512$ to $1024$ does not noticeably affect training speed, but it does require more memory. Nevertheless, this trade-off appears justified given the improved malware classification results. Overall, these results highlight the potential of leveraging language models associated with very complete reports on malware for automated classification. It is thus important to create such reports, gathering as much information as possible, from static and behavioral analysis.
2.4 Multi-CNN classification
To benchmark our results, we use a multi-convolutional neural network defined in [23], where we used Grayscale’s method [19] to transform the different sections of the PE files in images, and then train one CNN per section before gathering all the results thanks to a scoring function. Each CNN contains five layers, and is optimized with Adam [13], with a batch size of 32. This model is fast to train and the preprocessing is now a very common while intuitive tool to use, so it is interesting to see if this “naive” model has already good performances or if we indeed need to work with a more elaborated model.
Here we present the confusion matrix in Figure 11 and the classification metrics in Table 8 only for the test dataset.
Figure 11: Confusion matrix for the test dataset wit the multi-CNN.
Table 8: Testing classification report with multi-CNN
Class Accuracy Recall F1-score Support
backdoor 0.49 0.99 0.66 729
downloader 0.00 0.00 0.00 103
dropper 0.00 0.00 0.00 72
infostealer 0.00 0.00 0.00 44
ransomware 0.00 0.00 0.00 82
trojan 0.95 0.81 0.87 2940
virus 0.00 0.00 0.00 19
worm 0.93 0.96 0.95 1651
Global
Accuracy 0.83 5640
Macro avg 0.30 0.34 0.31 5640
Weighted avg 0.83 0.83 0.82 5640
We can see in Figure 11 that all the under-represented categories are labeled as backdoors, showing the incapacity of the model to learn correctly on an unbalanced dataset. In Table 8, the accuracy is decent only for trojan and worms, which are the two most represented classes of the dataset.
2.5 Balanced Dataset
In order to complete our tests, we run our classification task on a balanced dataset. We isolate the five more represented classes - namely trojans, worms, backdoors, downloaders and ransomwares - and take 813 samples from each of them. As usual, we use a train/test split of 80/20. We only test ModernBERT on this dataset, because here we just want to show that the previous unbalanced dataset does not bias the accuracy results. Results are illustrated in Table 9 for the train set and in Table 10 for the test set. We also provide the confusion matrices in Figures 12 and 13.
As we can see in Table 10, the smallest accuracy is 0.80, and concerns the downloader class. These results show that even with a balanced dataset our semantic preprocessing is well-adapted to a classification task using LLMs.
Table 9: Training classification report on balanced dataset with ModernBERT
Classe Accuracy Recall F1-score Support
backdoor 0.98 0.99 0.98 633
downloader 0.99 0.99 0.99 653
ransomware 1.00 1.00 1.00 641
trojan 0.97 0.92 0.94 671
worm 0.95 0.98 0.97 654
Global
Accuracy 0.98 3252
Macro avg 0.98 0.98 0.98 3252
Weighted avg 0.98 0.98 0.98 3252
Table 10: Testing classification report on balanced dataset with ModernBERT
Classe Accuracy Recall F1-score Support
backdoor 0.95 0.93 0.94 180
downloader 0.92 0.92 0.92 160
ransomware 0.93 0.98 0.95 172
trojan 0.80 0.79 0.79 142
worm 0.97 0.97 0.97 159
Global
Accuracy 0.92 813
Macro avg 0.92 0.92 0.92 813
Weighted avg 0.92 0.92 0.92 813
Figure 12: Confusion matrix for the balanced train dataset with ModernBERT.
Figure 13: Confusion matrix for the balanced test dataset with ModernBERT.